Solve the partial differential equation: r+ (a+b)s + abt = xy
1
Expert's answer
2020-10-06T18:53:30-0400

Solve by Monge’s method

Comparing the given equation with "Rr+Ss+Tt=V,"we have "R=1, S=a+b,T=ab,V=xy."

Here Monge’s subsidiary equations


"Rdpdy+Tdqdx-Vdydx=0"

"Rdy^2-Sdxdy+Tdx^2=0"


become


"dpdy+abdqdx-xydydx=0 \ \ \ \ \ \ \ \ \ \ (1)"

"dy^2-(a+b)dxdy+abdx^2=0 \ \ \ \ \ \ \ \ (2)"

Equation "(2)" may be factorised as


"(dy-bdx)=0 \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ (3)"

"(dy-adx)=0 \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ (4)"


On integration


"y-bx=c_1 \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ (5)"

"y-ax=c_2 \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ (6)"

Combining equation (3) with subsidiary equation (1), we get


"dp(bdx)+abdqdx-xy(bdx)dx=0"

"dp+adq-xydx=0"

By using (5)


"dp+adq-x(c_1+bx)dx=0"

On integration, we get


"p+aq-(\dfrac{c_1}{2})x^2-(\dfrac{b}{3})x^3=c_3"

By using (5)


"p+aq-\dfrac{x^2}{2}(y-bx)-(\dfrac{b}{3})x^3=c_3"

"p+aq-(\dfrac{1}{2})x^2y+(\dfrac{1}{6})bx^3=c_3"

Therefore the first intermediate integral is


"p+aq-\dfrac{1}{2}x^2y+\dfrac{1}{6}bx^3=f_1(y-bx)\ \ \ \ \ \ \ \ (7)"

Similarly, the second intermediate integral corresponding to equation (4) is


"p+bq-\dfrac{1}{2}x^2y+\dfrac{1}{6}ax^3=f_2(y-ax)\ \ \ \ \ \ \ \ (8)"

Now from above two intermediate integrals (7) and (8), we deduce the values of p and q as


"q=\dfrac{1}{6}x^3+\dfrac{1}{a-b}\big[f_1(y-bx)-f_2(y-ax)\big]"

"p=\dfrac{1}{2}x^2y-\dfrac{1}{6}(a+b)x^3+\dfrac{1}{a-b}\big[af_2(y-ax)-bf_1(y-bx)\big]"

Substituting these values of "p" and "q" in "dz=pdx+qdy," we get


"dz=\dfrac{1}{2}x^2ydx-\dfrac{1}{6}(a+b)x^3dx+"

"+\dfrac{1}{a-b}\big[af_2(y-ax)dx-bf_1(y-bx)dx\big]+"

"+\dfrac{1}{6}x^3dy+\dfrac{1}{a-b}\big[f_1(y-bx)dy-f_2(y-ax)dy\big]"

Or


"dz=\dfrac{1}{6}(3x^2ydx+x^3dy)-\dfrac{1}{6}(a+b)x^3dx-"

"-\dfrac{1}{b-a}\big[af_2(y-ax)dx-bf_1(y-bx)dx\big]-"

"-\dfrac{1}{b-a}\big[f_1(y-bx)dy-f_2(y-ax)dy\big]"

Or


"dz=\dfrac{1}{6}d(x^3y)-\dfrac{1}{6}(a+b)x^3dx+"

"+\dfrac{1}{b-a}f_2(y-ax)(dy-adx)-"

"-\dfrac{1}{b-a}f_1(y-bx)(dy-bdx)"

Integrating, we get the required solution as


"z=\dfrac{1}{6}x^3y-\dfrac{1}{24}(a+b)x^4+\phi(y-ax)+\phi_1(y-bx)"